Solve the system by substitution. 3x – y = –7
x + y = –9

Answer by asha(30)

3x-y =-7 eqn.1
x +y =-9 eqn.2
from eqn. we get y =-9-x
plugging the value of y in eqn.1 we get
3x-(-9-x)=-7
3x+9+x=-7
4x =-7-9 =-16
x=-16/4 =-4
but y =-9-x =-9-(-4)
y=-9+4 =-5
the answers are x=-4 and y=-5
good luck !!!